# Models That Prove Their Own Correctness

### v2

@@ 4.2 Reinforcement Learning from Verifier Feedback (RLVF)
We prove that RLVF learner can estimate the Verifiability gradient of $P_{\theta}$ using emulation alone in Lemma A.6. From a broader perspective, RLVF can be derived by viewing Self-Proving as a reinforcement learning problem in which the agent (prover) is rewarded when the verifier accepts. Indeed, RLVF is the Policy Gradient method (Sutton et al., 1999) for a verifier-induced reward. Convergence bounds for Policy Gradient methods are a challenging and active area of research (e.g. Agarwal et al. 2021), and so we leave the full analysis to future work.
@@ 4.3 Learning from annotated transcripts
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We adapt Sections 3 and 4 to this setting in Appendix C, where we present Annotated Transcripts. This can be viewed as adding Chain-of-Thought (Wei et al., 2022) to the model. The Transcript Learning algorithm naturally extends to annotated transcripts as well.
@@ 5.1 Setup: Training transformers to predict the GCD of two integers
We choose the input distribution $\mu$ to be the log-uniform distribution on $[10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton (2024), to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (Vaswani et al., 2017) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Appendix E.
@@
Following Charton (2024) as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Appendix D. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
@@
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Appendix D, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
  • Soundness: If $y\neq GCD(x_{1},x_{2})$, then $V(\mathbf{x_{1}x_{2}y\pi})$ rejects for any alleged proof $\mathbf{\pi}\in\Sigma^{\ast}$.
@@ 5.2 Models generalize beyond annotations
Figure 2: Verifiability with increasing amounts of annotation. $T$ is the number of steps added in Annotated Transcript Learning. Dashed lines indicate Euclidean depth, that bound the Verifiability of models that prove only for integers up to a certain number of steps. Each $T$ was run with three seeds, with mean $\pm$ standard error depicted.
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Appendix D; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
Figure 2 shows how $T$ affects the Verifiability of the learned model. As suggested by Lee et al. (2024), training the model on more intermediate steps results in better performance; in our case, increasing the number of intermediate steps $T$ yields better Self-Proving models. One might suspect that models only learn to execute the Euclidean algorithm in-context. To rule out this hypothesis, we derive an upper bound on the possible efficacy of such limited models. This bound is based on the Euclidean depth of integers $(x_{1},x_{2})$, which we define as the number of intermediate steps that the Euclidean algorithm makes before terminating on input $(x_{1},x_{2})$. Indeed, a model that only learns the to compute (in-context) the simple arithmetic of the Euclidean algorithm would only be able to prove the correctness of inputs $(x_{1},x_{2})$ whose depth does not exceed the annotation cutoff $T$.
@@ Acknowledgments
We are grateful to Micah Carroll and Avishay Tal for their helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
Appendix A Theoretical analyses for Section 4
@@
In this section we provide a formal description and analysis of Transcript Learning (TL, Section 4.1) and Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2). In Section A.1 we prove a convergence theorem for TL under convexity and Lipschitzness assumptions. Obtaining an analogous result for RLVF is more challenging; in lieu of a full analysis, we provide a lemma showing that the gradients estimated in the algorithm approximate the Verifiability of the model in Section A.2.
@@
Definition A.1 (Self-Proving model learner).
@@
  • A family of autoregressive models $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$ where $d\in\mathbb{N}$ is the number of parameters in the family. Recall (Section 4) that for each $\theta$ and $z\in\Sigma^{\ast}$, the random variable $P_{\theta}(z)$ is determined by the logits $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. For any $z\in\Sigma^{\ast}$ and $\sigma\in\Sigma$, the learner $\Lambda$ can compute the gradient of the $\sigma^{\text{th}}$ logit, that is, $\nabla_{\theta}\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$.
  • Sample access to a the input distribution $\mu$. That is, $\Lambda$ can sample $x\sim\mu$.
@@
We remark that analysis of Transcript Learning will require a slight strengthening of the final item above. This is discussed in Section A.1.
Throughout this section, we will refer to the transcript of an interaction between a verifier and a prover (see Figure 1). We will this transcript by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$, and for any index $s\in|\pi|$ we will write $\pi_{<s}\in\Sigma^{s-1}$ to denote the $s$-long prefix of $\pi$.
A.1 Transcript Learning
@@
Definition A.2 (Query generator).
@@
We assume that access to the verifier specification (Definition A.1) includes access to the query generator. After all, the verifier—who is assumed to be efficient—samples from $V_{q}$ during the interaction. Moreover, we will assume that for any partial interaction and any sequence $q^{\prime}$, the learner is able to compute the probability that $q^{\prime}$ was the next query. In other words, we assume the learner can compute the probability density function of $V_{q}$.
A transcript generator is a random variable over transcripts that faithfully represents the interaction of the verifier with some prover for a given input. An honest transcript generator is one who is fully supported on transcripts accepted by the verifier.
Definition A.3 (Honest transcript generator).
@@
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition A.3 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition 3.2). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning (Algorithm 1 described below).
@@
1:Initialize $\theta_{0}\coloneqq\vec{0}$.
22:for $i=0,\dots,N-1$ do
3:Sample $x\sim\mu$ and $\pi^{*}=(y,q_{1},a_{1},\dots,q_{R},a_{R})\sim\mathcal{T}_{V}^{\ast}(x)$. Denote $a_{0}\coloneqq y$.
3      4:foreach Round of interaction $r=0,\dots,R$ do
4            5:Let $S(r)$ denote the indices of the $r^{\text{th}}$ answer $a_{r}$ in $\pi^{\ast}$. for $s\in S(r)$ do
5                  6:Compute # Forwards and backwards pass
$\alpha_{s}(\theta_{i})$ $\coloneqq\Pr_{\sigma\sim p_{\theta_{i}}(x\pi_{<s})}[\sigma=\pi_{s}]$
$\vec{d}_{s}(\theta_{i})$ $\coloneqq\nabla_{\theta}\alpha_{s}(\theta_{i})=\nabla_{\theta}\log\Pr_{\sigma\sim p_{\theta_{i}}(x\pi_{<s})}[\sigma=\pi_{s}].$
7:If $r\geq 1$, let $q_{r}$ denote the $r^{\text{th}}$ query $q_{r}$ in $\pi^{\ast}$, and let $t$ denote its first index. That is, $\pi_{<t}^{\ast}=(y,q_{1},a_{1},\dots,q_{t-1},a_{t-1})$. Compute # Emulate the verifier
$\beta_{r}(\theta_{i})\coloneqq\Pr_{q^{\prime}\sim V_{q}(x\pi^{\ast}_{<t})}[q^{\prime}=q].$
6            8:Update
$\theta_{i+1}\coloneqq\theta_{i}+\lambda\cdot\alpha_{0}(\theta_{i})\cdot\left(\prod_{\begin{subarray}{c}r\in[R]\\ s\in S(r)\end{subarray}}\beta_{r}(\theta_{i})\alpha_{s}(\theta_{i})\right)\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i}).$
9:Output $\bar{\theta}\coloneqq\frac{1}{N}\sum_{i\in[N]}\theta_{i}$.

### v3

@@ 4.2 Reinforcement Learning from Verifier Feedback (RLVF)
We prove that RLVF learner can estimate the Verifiability gradient of $P_{\theta}$ using emulation alone in Lemma B.6. From a broader perspective, RLVF can be derived by viewing Self-Proving as a reinforcement learning problem in which the agent (prover) is rewarded when the verifier accepts. Indeed, RLVF is the Policy Gradient method (Sutton et al., 1999) for a verifier-induced reward. Convergence bounds for Policy Gradient methods are a challenging and active area of research (e.g. Agarwal et al. 2021), and so we leave the full analysis to future work.
@@ 4.3 Learning from annotated transcripts
However, when training Self-Proving model, it may be useful for it to first generate an “annotated” answer $\widetilde{a}$ which is then trimmed down to the actual answer $a$ to be sent to the verifier. We adapt Sections 3 and 4 to this setting in Appendix D, where we present Annotated Transcripts. The TL and RLVF algorithms naturally extend to annotated transcripts as well. Table 2 shows that annotations significantly improve performance of TL.
Annotations can be viewed as adding Chain-of-Thought (Wei et al., 2022). As a concrete example, consider our experiments on computing the GCD. As detailed in Section 5.2, a proof $\pi$ in this setting is the output of an iterative process—the extended Euclidean algorithm—starting from the input $x$: $x\mapsto\pi_{1}\mapsto\pi_{2}\mapsto\dots\mapsto\pi$. The annotation of the proof $\pi$ consists the first $T$ steps $(\pi_{1},\dots,\pi_{T})$ up to some fixed cutoff $T$. These are prepended to the proof and shown to the model during TL training. At inference time, the model is evaluated only on whether it generated the proof $\pi$ correctly.
@@ 5.1 Setup: Training transformers to predict the GCD of two integers
We choose the input distribution $\mu$ to be the log-uniform distribution on $[10^{4}]$, and train the transformer on sequences of the form $\mathbf{(x_{1}x_{2}y)}$, where $x_{1},x_{2}\sim\mu$ and $y=GCD(x_{1},x_{2})$. This is a scaling-down of Charton (2024), to allow single GPU training of Self-Proving transformers. In all of our experiments, we use a GPT model (Vaswani et al., 2017) with 6.3M parameters trained on a dataset of 1024K samples in batches of 1024. Full details are deferred to Appendix F.
@@
Following Charton (2024) as a baseline, we find that transformers can correctly compute the GCD with over $99\%$ probability over $(x_{1},x_{2})\sim\mu$. To what extent can they prove their answer? To answer this question, we first devise a natural proof system based on Bézout’s theorem. Its specification and formal guarantees are deferred to Appendix E. We denote its verification algorithm by $V$, and highlight some important features of the experimental setup:
@@
  • Completeness: For any $x_{1},x_{2},y\in[10^{4}]$ such that $y=GCD(x_{1},x_{2})$, there exists a proof $\mathbf{\pi}$ such that $V(\mathbf{x_{1}x_{2}y\pi})$ accepts. As detailed in Appendix E, the proof $\pi$ consists of a pair of integers who are Bézout coefficients for $x_{1},x_{2}$.
  • Soundness: If $y\neq GCD(x_{1},x_{2})$, then $V(\mathbf{x_{1}x_{2}y\pi})$ rejects [^6] With probability 1, i.e., $s=0$ in Definition 3.2. for any alleged proof $\mathbf{\pi}\in\Sigma^{\ast}$.
@@
Table 2 shows our main experimental result, which has the following key takeaways:
  1. 1.
@@
  2. 2.
    A base Self-Proving Model with fairly low Verifiability of 40% can be improved to 79.3% via Reinforcement Learning from Verifier Feedback (RLVF). Although it does not rely on honest transcripts, RLVF trains slowly: this nearly-twofold improvement took four million iterations.
  3. 3.
    Most efficient is Annotated Transcript Learning, which yielded a model with 96% Verifiability in 100K iterations.
We further investigate the effect of annotations next.
@@ 5.2 Models generalize beyond annotations
Figure 2: Verifiability with increasing amounts of annotation. $T$ is the number of steps added in Annotated Transcript Learning. Dashed lines indicate Euclidean depth, that bound the Verifiability of models that prove only for integers up to a certain number of steps. Each $T$ was run with three seeds, with mean $\pm$ standard error depicted. The upper graph provides a zoomed-in view of the 82% to 98% range from the lower graph, which spans a broader scale from 20% to 100%.
The proof $\mathbf{\pi}$ is annotated by including intermediate steps in its computation. Details are deferred to Appendix E; roughly speaking, we observe that the proof $\mathbf{\pi}$ for input $\mathbf{(a,b)}$ is obtained as the last element in a sequence $\mathbf{a,b,\pi_{1},\pi_{2},\dots}$ computed by the Euclidean algorithm. We annotate the proof $\mathbf{\pi}$ by prepending to it the sequence of Euclidean steps $\mathbf{(\pi_{1},\dots,\pi_{T})}$ up to some fixed cutoff $T$.
Figure 2 shows how $T$ affects the Verifiability of the learned model. As suggested by Lee et al. (2024), training the model on more intermediate steps results in better performance; in our case, increasing the number of intermediate steps $T$ yields better Self-Proving models. One might suspect that models only learn to execute the Euclidean algorithm in-context. To rule out this hypothesis, we derive an upper bound on the possible efficacy of such limited models. This bound is based on the Euclidean depth of integers $(x_{1},x_{2})$, which we define as the number of intermediate steps that the Euclidean algorithm makes before terminating on input $(x_{1},x_{2})$. Indeed, a model that only learns to compute (in-context) the simple arithmetic of the Euclidean algorithm would only be able to prove the correctness of inputs $(x_{1},x_{2})$ whose depth does not exceed the annotation cutoff $T$.
@@ 5.3 Base of representation
6 Limitations
Our experiments are focused on a single ground-truth capability, namely, computing the GCD. Yet, the theoretical portion of our work holds for any ground-truth $F^{\ast}$ that admits an Interactive Proof system. Training large Self-Proving models for more complex ground-truths will likely pose additional practical learning challenges. With that said, we stress that generating accepting transcripts for use in Transcript Learning is distinct from these learning challenges. Collecting accepting transcripts is a purely computational task, and can even be done “offline” prior to the model’s training.
Additionally, in our current learning methods, each individual ground-truth capability requires training a separate Self-Proving model. It would be interesting to adapt our definition and methods to deal with a single generalist Self-Proving model that proves its correctness to multiple verifiers of different ground-truths.
@@ Acknowledgments
We are grateful to Micah Carroll, Mojtaba Yaghoobzadeh, Avishay Tal, and anonymous reviewers for their helpful comments. This research was supported by DARPA-TA1 under grant no. HR001119S0076, and by the Simons Collaboration on the Theory of Algorithmic Fairness.
Appendix A A definition for general loss functions and one-to-many relations
We present a variant of Self-Proving models (Definition 3.4) generalized in two ways.
General (bounded) loss functions.
In Definition 3.1 we implicitly use the 0-1 loss when measuring the correctness of a model: For any $x\in X$, we measure only whether the model generated the correct output $y=F^{\ast}(x)$, but not how “far” the generated $y$ was from $F^{\ast}(x)$. It is often the case in machine learning that we would be satisfied with models that generate a “nearly-correct” output. This is formalized by specifying a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$ and measuring the probability that $\ell(x,y)$ is smaller than some threshold $\lambda\in[0,1)$, where $x$ is drawn from the input distribution $\mu$, and $y$ is generated by the model when given input $x$.
In the context of language modeling, different loss function allow for a more fine-grained treatment of the semantics of a given task. As an example, consider the prime-counting task:
  • Given an integer $x<10^{9}$, output the number of primes less than or equal to $x$.
In the notation of Section 3, the prime-counting task would be captured by the ground-truth function
$F^{\ast}(x)\coloneqq\left|\left\{p\in\mathbb{N}\ \middle|\ p\leq x,\ p\textrm{is prime}\right\}\right|.$
Per Definition 3.1, any output other than $F^{\ast}(x)$ is “just as incorrect” as any other. Yet, we might prefer outputs that are closer to the correct answer, say, in $L_{1}$ norm. This preference can be captured by the following bounded loss function
$\ell_{1}(x,y)\coloneqq\begin{cases}|y-F^{\ast}(x)|\cdot 10^{-9}\quad&\textrm{if }y\leq 10^{9}\\ 1&\textrm{else.}\end{cases}$
In particular, if we are interested in knowing the answer only up to some additive constant $C$, we could say that an output $y$ is “correct-enough” if $\ell_{1}(x,y)\leq C\cdot 10^{-9}$.
More generally, we relax Definition 3.1 to capture approximate correctness as follows.
Definition A.1 (Approximate correctness).
Let $\mu$ be a distribution over input sequences in $\Sigma^{\ast}$ and let $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$ be a loss function. For any $\alpha,\lambda\in[0,1]$, we say that model $F_{\theta}$ is $(\alpha,\lambda)$-correct with respect to $\mu$ if
$\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim F_{\theta}(x)\end{subarray}}[\ell(x,y)\leq\lambda]\geq\alpha.$
One-to-many-relations.
In Section 3, we focused on the setting of models of a ground-truth function $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$. That is, when each input $x$ has exactly one correct output, namely $F^{\ast}(x)$. A more general setting would be to consider a ground-truth relation $L\subseteq\Sigma^{\ast}\times\Sigma^{\ast}$. Then, we say that $y$ is a correct output for $x$ if $(x,y)\in L$. Importantly, this allows a single $x$ to have many possible correct outputs, or none at all.
Note that we must take care to choose a loss function $\ell$ that captures correctness with respect to the relation $L$, i.e., $\ell(x,y)=0$ if and only if $(x,y)\in L$. Equivalently, any loss function $\ell$ induces a relation $L\coloneqq\{(x,y)\ |\ \ell(x,y)=0\}$. Therefore, our relaxation to approximate-correctness Definition A.1 already captures the setting of one-to-many relations, since an input $x$ may have multiple $y^{\ast}$ such that $\ell(x,y^{\ast})=0$.
A.1 The general definition
We first present a relaxed definition of Interactive Proof systems for verifying approximate-correctness.
Definition A.2 (Definition 3.2, generalized).
Fix a soundness error $s\in(0,1)$, a threshold $\lambda\in[0,1)$, a finite set of tokens $\Sigma$, and a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$. A verifier $V$ for $\ell$ with threshold $\lambda$ is a probabilistic polynomial-time algorithm that is given explicit inputs $x,y\in\Sigma^{\ast}$ and black-box (oracle) query access to a prover $P$. It interacts with $P$ over $R$ rounds (see Figure 1) and outputs a decision $V,P\right>(x,y)\in\{\mathrm{reject},\mathrm{accept}\}$. Verifier $V$ satisfies the following two guarantees:
  • Completeness: There exists an honest prover $P^{\ast}$ such that, for all $x,y\in\Sigma^{\ast}$, if $\ell(x,y)=0$ then
    $\Pr[\left<V,P^{\ast}\right>\!(x,y)\ \mathrm{accepts}]=1,$
    where the probability is over the randomness of $V$.
  • Soundness: For all $P$ and for all $x,y\in\Sigma^{\ast}$, if $\ell(x,y)>\lambda$ then
    $\Pr[\left<V,P\right>(x,y)\ \mathrm{accepts}]\leq s,$
Indeed, for a given ground-truth function $F^{\ast}\colon\Sigma^{\ast}\to\Sigma^{\ast}$, Definition 3.2 can be recovered by choosing the 0-1 loss
$\ell_{F^{\ast}}(x,y)\coloneqq\begin{cases}1\quad\mathrm{if}\ x\neq F^{\ast}(y)\\ 0\quad\mathrm{else}.\end{cases}$
and any threshold $\lambda\in[0,1)$.
Remark A.3 (Connection to Interactive Proofs of Proximity).
Definition A.2 can be seen as a slight generalization of (perfect completeness) Interactive Proofs of Proximity (IPPs, Rothblum et al. 2013). An IPP for a relation $L\subseteq\Sigma^{\ast}\times\Sigma^{\ast}$ with proximity parameter $\lambda$ is obtained by instantiating Definition A.2 with the loss function $\ell_{\mathrm{Hamming}}$ defined by
$\ell_{\mathrm{Hamming}}(x,y)\coloneqq\min\left\{\frac{\#\{i\ |\ y_{i}\neq y^{\ast}_{i}\}}{|y|}\ \middle|\ (x,y^{\ast})\in L,\ |y^{\ast}|=|y|\right\},$
that is, $\ell_{\mathrm{Hamming}}(x,y)$ is the fraction of tokens in $y$ that must be changed so as obtain an output $y^{\ast}$ with $(x,y^{\ast})\in L$. However, the motivation of Rothblum et al. [2013] was studying sublinear time verification, whereas ours is to relax the requirements of traditional Interactive Proofs towards meeting common desiderata in machine learning.
With this relaxed notion of Interactive Proofs in hand, we are now ready to define Self-Proving models for general (bounded) loss functions.
Definition A.4 (Definition 3.4, generalized).
Fix a loss function $\ell\colon\Sigma^{\ast}\times\Sigma^{\ast}\to[0,1]$, a verifier $V$ for $\ell$ with threshold $\lambda\in[0,1)$ as in Definition A.2, and a distribution $\mu$ over inputs $\Sigma^{\ast}$. The Verifiability of a model $P_{\theta}\coloneq\Sigma^{\ast}\to\Sigma^{\ast}$ is defined as
$\mathrm{ver}_{V,\mu}(\theta)\coloneqq\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[\left<V,P_{\theta}\right>(x,y)\ \mathrm{accepts}\right].$
Analogously to Remark 3.5, we observe that Verifiability (Definition A.4) implies approximate-correctness: Suppose $P_{\theta}$ is $\beta$-Self-Proving model with respect to a verifier $V$ that has soundness error $s$ and threshold parameter $\lambda$ for loss function $\ell$. Then by a union bound,
$\Pr_{\begin{subarray}{c}x\sim\mu\\ y\sim P_{\theta}(x)\end{subarray}}\left[\ell(x,y)\leq\lambda\right]\geq\beta-s.$
Importantly, as emphasized throughout this paper, soundness of $V$ implies that for all inputs $x$, any output $y$ such that $\ell(x,y)>\lambda$ is rejected with high probability ($1-s$).
Appendix B Theoretical analyses for Section 4
@@
In this section we provide a formal description and analysis of Transcript Learning (TL, Section 4.1) and Reinforcement Learning from Verifier Feedback (RLVF, Section 4.2). In Section B.1 we prove a convergence theorem for TL under convexity and Lipschitzness assumptions. Obtaining an analogous result for RLVF is more challenging; in lieu of a full analysis, we provide a lemma showing that the gradients estimated in the algorithm approximate the Verifiability of the model in Section B.2.
@@
Definition B.1 (Self-Proving model learner).
@@
  • A family of autoregressive models $\{P_{\theta}\}_{\theta\in\mathbb{R}^{d}}$ where $d\in\mathbb{N}$ is the number of parameters in the family. Recall (Section 4) that for each $\theta$ and $z\in\Sigma^{\ast}$, the random variable $P_{\theta}(z)$ is determined by the logits $\log p_{\theta}(z)\in\mathbb{R}^{|\Sigma|}$. For any $z\in\Sigma^{\ast}$ and $\sigma\in\Sigma$, the learner $\Lambda$ can compute the gradient of the $\sigma^{\text{th}}$ logit, that is, $\nabla_{\theta}\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$. In particular, $\log\Pr_{\sigma^{\prime}\sim p_{\theta}(z)}[\sigma=\sigma^{\prime}]$ is always differentiable in $\theta$.
  • Sample access to the input distribution $\mu$. That is, $\Lambda$ can sample $x\sim\mu$.
@@
Throughout this section, we will refer to the transcript of an interaction between a verifier and a prover (see Figure 1). We will denote this transcript by $\pi=(y,q_{1},a_{1},\dots,q_{R},a_{R})$, and for any index $s\in[|\pi|]$ we will write $\pi_{<s}\in\Sigma^{s-1}$ to denote the $s$-long prefix of $\pi$.
B.1 Transcript Learning
@@
Definition B.2 (Query generator).
@@
A transcript generator is a random variable over transcripts that faithfully represents the interaction of the verifier with some prover for a given input. An honest transcript generator is one who is fully supported on transcripts accepted by the verifier. We denote accepting transcripts by $\pi^{*}=(y^{*},q^{*}_{1},a^{*}_{1},\dots,q^{*}_{R},a^{*}_{R})$.
Definition B.3 (Transcript generator).
@@
Notice that for any verifier $V$, there is a one-to-one correspondence between transcript generators and (possibly randomized) provers. We intentionally chose not to specify a prover in Definition B.3 to emphasize that transcripts can be “collected” independently of the honest prover (see completeness in Definition 3.2), and in fact can be collected “in advance” prior to learning (see Figure 4). As long as the generator is fully supported on honest transcripts, it can be used for Transcript Learning (Algorithm 1 described next).
@@
2Initialize $\theta_{0}\coloneqq\vec{0}$.
3 for $i=0,\dots,N-1$ do
4       Sample $x\sim\mu$ and $\pi^{*}=(y^{*},q^{*}_{1},a^{*}_{1},\dots,q^{*}_{R},a^{*}_{R})\sim\mathcal{T}_{V}^{\ast}(x)$. Denote $a_{0}\coloneqq y^{*}$.
5       foreach Round of interaction $r=0,\dots,R$ do
6             Let $S(r)$ denote the indices of the $r^{\text{th}}$ answer $a_{r}$ in $\pi^{\ast}$, and let $\pi_{<s}$ denote the prefix of the partial transcript $(y,q^{*}_{1},a_{1}^{*},\dots,q^{*}_{r})$.
7             for $s\in S(r)$ do
8                   Compute # Forwards and backwards pass
$\alpha_{s}(\theta_{i})$ $\coloneqq\Pr_{\sigma\sim p_{\theta_{i}}(x\pi_{<s})}[\sigma=\pi^{*}_{s}]$
$\vec{d}_{s}(\theta_{i})$ $\coloneqq\nabla_{\theta}\log\alpha_{s}(\theta_{i})=\nabla_{\theta}\log\Pr_{\sigma\sim p_{\theta_{i}}(x\pi_{<s})}[\sigma=\pi^{*}_{s}].$
9      Update
$\theta_{i+1}\coloneqq\theta_{i}+\lambda\cdot\prod_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\alpha_{s}(\theta_{i})\cdot\sum_{\begin{subarray}{c}r\in[R]\cup\{0\}\\ s\in S(r)\end{subarray}}\vec{d}_{s}(\theta_{i}).$
Output $\bar{\theta}\coloneqq\frac{1}{N}\sum_{i\in[N]}\theta_{i}$.